Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario.
Part 2:
Solve the system and explain what the results mean according to your scenario.
2x + y = 11
3x + 4y = 24

Answer by lenny460(1073)

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario.
Part 2:
Solve the system and explain what the results mean according to your scenario.
2x + y = 11
3x + 4y = 24


2x + y = 11
2x + y - 2x = 11 - 2x
2x - 2x = 0
Therefore:
y = 11 - 2x


Substitute y = 11 - 2x into equation (2)


3x + 4y = 24
3x + 4(11 - 2x) = 24
3x + 44 - 8x = 24
3x - 8x = -5x
-5x + 44 = 24
-5x + 44 - 44 = 24 - 44
44 - 44 = 0
24 - 44 = -20
-5x = -20
x = 4


Substitute x = 4 into equation (1).

2x + y = 11
2(4) + y = 11
8 + y = 11
8 + y - 8 = 11 - 8
8 - 8 = 0
11 - 8 = 3
y = 3



Answer:
x = 4 and
y = 3






Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com